# Facial Expression Analysis and Its Potentials in IoT Systems: A Contemporary Survey

### v1

@@ Abstract.
Facial expressions convey human emotions and can be categorized into macro-expressions (MaEs) and micro-expressions (MiEs) based on duration and intensity. While MaEs are voluntary and easily recognized, MiEs are involuntary, rapid, and can reveal concealed emotions. The integration of facial expression analysis with Internet-of-Thing (IoT) systems has significant potential across diverse scenarios. IoT-enhanced MaE analysis enables real-time monitoring of patient emotions, facilitating improved mental health care in smart healthcare. Similarly, IoT-based MiE detection enhances surveillance accuracy and threat detection in smart security. This work aims at providing a comprehensive overview of research progress in facial expression analysis and explores its integration with IoT systems. We discuss the distinctions between our work and existing surveys, elaborate on advancements in MaE and MiE techniques across various learning paradigms, and examine their potential applications in IoT. We highlight challenges and future directions for the convergence of facial expression-based technologies and IoT systems, aiming to foster innovation in this domain. By presenting recent developments and practical applications, this study offers a systematic understanding of how facial expression analysis can enhance IoT systems in healthcare, security, and beyond.
Facial expression analysis, Internet of Thing, macro- and micro-expressions
Nomenclature
Abbreviations Definitions
AFEW 7.0 The acted facial expressions in the wild
AU Action unit
BERT Bidirectional encoder representations from transformers
BU-3DFE Binghamton university 3D facial expression
CAER Context-aware emotion recognition
CASME Chinese Academy of Sciences micro-expression
CASME II Chinese Academy of Sciences micro-expression II
CAS(ME)2 Chinese Academy of Sciences macro-expressions and micro-expressions
CK+ The Extended Cohn-Kanade Dataset
CNN Convolutional neural network
DFEW Dynamic facial expression in the wild
ExpW Expression in-the-wild
FACS Facial action coding system
FER-2013 Facial expression recognition 2013
GAN Generative adversarial network
HOG Histogram of gradient
ISED Indian spontaneous expression database
IoT Internet of Things
JAFFE Japanese female facial expression
KDEF Karolinska directed emotional faces
LBP Local binary pattern
LBP-TOP Local binary pattern from three orthogonal planes
LSTM Long short-term memory
MaE Macro-expression
MDMO Main directional mean optical-flow
MEVIEW Micro-expression videos in the wild
MiE Micro-expression
MMEW Micro-and-macro in expression warehouse
PEDFE Padova emotional dataset of facial expressions
RaFD Radboud faces database
RNN Recurrent neural network
ROI Region of interest
SAMM Spontaneous actions and micro-movements
SFEW 2.0 Static facial expressions in the wild
SMIC Spontaneous micro-expression corpus
TFD The Toronto face database
UAR Mean class accuracy
ViT Vision transformer
WAR Overall accuracy
@@ 1. Introduction
Since its inception in 1999, Internet-of-Things (IoT) technology has seamlessly integrated into the fabric of modern society, driving key innovations in both civilian and industrial sectors. The widespread adoption of IoT technology generates skyrocketing amounts of daily data that were traditionally processed in clusters of cloud servers. Due to the ever-increasing privacy concerns, the information process is shifted from the cloud to the network edge. The low access latency and high data security of edge computing enable new application domains, such as emotion detection, traffic surveillance, and healthcare (Dong et al., 2019, 2018).
Figure 1. A framework of edge-driven facial expression analysis.
Figure 2. An illustration of MaEs and MiEs  in micro-and-macro in expression warehouse (Ben et al., 2021).
@@
Since facial expressions convey 55% of emotional messages during the exchange of moods, thoughts, feelings, and mental states, they have become a medium for sensing emotions in front of facility screens (Mehrabian and Wiener, 1967). Recent advancements in IoT technologies have driven the widespread application of facial expression analysis in practical scenarios. As shown in Fig. 1, IoT devices can first collect raw facial expression data for pre-processing. To fully leverage low access latency and high data security during the implementation of real-time privacy-preserving systems, the pre-processed features are then uploaded to edge devices for decision-making. Facial expressions can be divided into macro-expressions (MaEs) and micro-expressions (MiEs) based on their duration and intensity. Two types of emotions within MaEs and MiEs are illustrated in Fig.2. In terms of duration, MaEs approximately last 0.5 to 4 seconds, while MiEs last less than 0.5 seconds (Shen et al., 2012). The short duration of MiEs conveys subtle emotional message, which makes emotion perception from MiEs a challenging task without specialized techniques and training. The major differences between MaEs and MiEs are described as follows.
  • MaEs can be perceived during regular interactions. The recognition accuracy of MaEs  can exceed 97% in a controlled laboratory due to the spontaneity and noise immunity. The MaEs can be applied in multiple scenarios, e.g., autonomous driving, psychological health assessment, and human-computer interaction.
@@
  • MiEs are involuntary and rapid. Due to the short duration and subtle emotional cues, the detection of MiEs  can be quite challenging with the naked eye. It is reported that the average recognition accuracy of MiEs is around 47% even after specialized training (Frank et al., 2009). Nevertheless, MiEs can reveal the true emotions of a person since they are instinctive and cannot be concealed. Hence, numerous applications of MiEs in IoT systems are being explored, such as lie detection, criminal identification, and security control.
@@
The objective of facial expression analysis lies in detecting the emotions of humans from MaEs and MiEs. The research works on MaE analysis are mainly based on deep models that detect human emotions by leveraging the isolated static images and sequential dynamic images of MaEs. More specifically, the deep static MaE analysis performs well for two scenarios: (1) when several isolated images are available; and (2) when real-time expression recognition is required. The deep dynamic MaE analysis can benefit from the comprehensive understanding of temporal dynamics of MaEs. Since the emotion recognition of MiEs is challenging due to the short duration and subtle emotional cues, the research works on holographic MiE analysis consist of two parts, i.e., MiE spotting and MiE recognition. The MiE spotting aims at detecting MiEs within a given video, and the MiE recognition then uses the detected MiEs for MiE classification within a set of MiE categories. Different from the MaE analysis, the MiE analysis can provide valuable insights into potentially hidden human emotions. In this work, we aim at providing a contemporary tutorial on facial expression analysis that ranges from the used datasets to the detailed technical methods. Compared with the previous surveys on facial expression analysis, our review covers a broader perspective on both MaE and MiE analysis. The emerging techniques on MaE and MiE analysis are discussed on deep MaE recognition and holographic MiE analysis. Besides, we also discuss the potential integrations of MaE and MiE analysis with the IoT systems in order to inform new directions of facial expression application.
As shown in Fig. 3, the remaining work is organized as follows. Section 2 introduces the recent surveys on MaE and MiE analysis and major differences from our work. Section 3 presents the datasets for MaE and MiE analysis. Section 4 discusses the deep MaE recognition. Section 5 discusses the holographic MiE recognition that includes spotting and recognition procedures. Section 6 discusses the applications of MaE and MiE analysis in IoT systems. Section 7 discusses the challenges and future directions. The conclusions are presented in Section 8.
Figure 3. The organization of the remaining work.
@@ 2.2. Recent Surveys on MiE Analysis
Previous surveys on MiE analysis have comprehensively examined the field from various perspectives, e.g., MiE analysis, dataset creation, and psychological foundations of MiE development. However, critical aspects such as advanced MiE analysis and practical MiE applications remain underexplored. To address the aforementioned gaps, our work delves into state-of-the-art MiE analysis in Section 5. Additionally, we explore the transformative potential of MiE applications in the IoT system in Section 6.2. We expect to provide a holistic perspective that can bridge foundational research with cutting-edge advancements and wild applicability.
3. Datasets for Facial Expression Analysis
3.1. MaE Datasets
The recent MaE datasets including Japanese female facial expression (JAFFE) (Lyons et al., 1998), Karolinska directed emotional faces (KDEF) (Goeleven et al., 2008), MMI (Valstar et al., 2010; Pantic et al., 2005), Binghamton university 3D facial expression (BU-3DFE) (Yin et al., 2006), the Toronto face database (TFD) (Susskind et al., 2010), The Extended Cohn-Kanade Dataset (CK+) (Lucey et al., 2010), Radboud faces database (RaFD) (Langner et al., 2010), MUG (Aifanti et al., 2010), Multi-PIE (Gross et al., 2010), static facial expressions in the wild (SFEW 2.0) (Dhall et al., 2011), Oulu-CASIA (Zhao et al., 2011), facial expression recognition 2013 (FER-2013) (Goodfellow et al., 2013), Indian spontaneous expression database (ISED) (Happy et al., 2017), EmotioNet (Fabian Benitez-Quiroz et al., 2016), wild affective faces database (RAF-DB) (Li et al., 2017a), AffectNet (Mollahosseini et al., 2019), the acted facial expressions in the wild (AFEW 7.0) (Dhall et al., 2017), expression in-the-wild (ExpW) (Zhang et al., 2018b), Aff-Wild2 (Kollias and Zafeiriou, 2019), context-aware emotion recognition (CAER) (Lee et al., 2019), dynamic facial expression in the wild (DFEW) (Jiang et al., 2020), FERV39k (Wang et al., 2022c), Padova emotional dataset of facial expressions (PEDFE) (Miolla et al., 2023), CalD3r + MenD3s (Ulrich et al., 2024). Table 1 illustrates the detailed information of the MaE datasets.
@@ 3.1. MaE Datasets
RaFD (Langner et al., 2010) 67 1 1608 images 256 × 256
@@
Oulu-CASIA (Zhao et al., 2011) 36 2
@@
EmotioNet (Fabian Benitez-Quiroz et al., 2016) N/A N/A 1,000,000 images Multiple
@@
AFEW 7.0 (Dhall et al., 2017) N/A N/A 1809 sequences Multiple BEs N/A N/A N/A Movie P & S 2017
@@
Aff-Wild2 (Kollias and Zafeiriou, 2019) 458 N/A 558 sequences Multiple
@@
FERV39k (Wang et al., 2022c) N/A N/A 38,935 sequences 3840 × 2160 BEs N/A N/A N/A Internet P & S 2021
@@
  • 1 BEs: Basic expressions (seven); N/A: Not applicable; P: Posed; S: Spontaneous;
  • 2 NAF: Onset frame, Apex frame, Offset frame, respectively.
@@ 3.2. MiE Datasets
In recent years, multiple datasets have been built for further development of MiE analysis. Initially, researchers collected the MiE databases (e.g., Polikvsky’s database (Polikovsky et al., 2009) and USF-HD (Shreve et al., 2011)) by asking the participants to pose or mimic facial expressions. These posed datasets supported the preliminary studies of MiE, but were not used for current analysis due to the properties of private datasets and not capturing the nature of MiE. Later, more spontaneous datasets, including the spontaneous micro-expression corpus (SMIC) (Li et al., 2013), Chinese Academy of Sciences micro-expression (CASME) (Yan et al., 2013), Chinese Academy of Sciences micro-expression II (CASME II) (Yan et al., 2014), Chinese Academy of Sciences macro-expressions and micro-expressions (CAS(ME)2) (Qu et al., 2017), spontaneous actions and micro-movements (SAMM) (Davison et al., 2016), micro-expression videos in the wild (MEVIEW) (Husák et al., 2017), CAS(ME)3 (Li et al., 2022b), MMEW (Ben et al., 2021), and 4D spontaneous micro expression database (4DME) (Li et al., 2022a), were utilized in MiE research. The detailed information on the spontaneous MiE datasets is provided in Table 2.
@@
Figure 4. Samples of current MiE datasets. These are as follows: (a) SMIC (Li et al., 2013), (b) CASME (Yan et al., 2013), (c) CASME II (Yan et al., 2014), (d) CAS(ME)2 (Qu et al., 2017), (e) SAMM (Davison et al., 2016), (f) MEVIEW (Husák et al., 2017), (g) CAS(ME)3 (Li et al., 2022b), (h) MMEW (Ben et al., 2021), and (i) 4DME (Li et al., 2022a).
@@
Pons et.al . (Pons and Masip, 2018b) 2018 IN+FA N/A SFEW 2.0 60.5% CNN LOSO
Saurav et al. (Saurav et al., 2022) 2022 IN+FA+CE N/A FER2013/RAF-DB /CK+ /SFEW 2.0 72.77% /86.07% /98.54% /60.5% CNN LOSO
Shao et al. (Shao and Qian, 2019) 2019 FD+FA+DA RES CK+/BU-3DFE/FER2013 86.50%/86.50%/71.14% CNN 10 fold
@@
Wadhawan et al. (Wadhawan and Gandhi, 2023) 2023 FD+CE N/A CK+/JAFFE/SFEW 2.0 97.31%/97.14%/44.50% CNN 10 fold+LOSO
@@
Atabansi et al. (Atabansi et al., 2021) 2021 FD+DA N/A Oulu-CASIA 98.11% CNN Hold out
Multi-task Learning Pons et al. (Pons and Masip, 2018a) 2018 DA RES SFEW 2.0 45.9% CNN Hold out
@@
Yu et al. (Yu and Xu, 2022) 2022 DA+FA RES+Atten RAF-DB/SFEW 2.0/CK+/Oulu-CASIA 90.36%/45.78%/98.33%87.32% CNN 10 fold+LOSO
Liu et al. (Liu et al., 2024) 2023 N/A RES+Trans AffectNet/Aff-Wild2 65.80%/64.13 CNN Hold out
Yu et al. (Yu et al., 2020a) 2020 FA+ROI N/A CK+/Oulu-CASIA 99.08%/90.40% GRU 10 fold
Xiao et al. (Xiao et al., 2023) 2023 FD N/A CK+/MMI/RAF-DB 9.07%/84.62%/87.52% CNN 5 fold+Hold out
Zhao et al. (Zhao et al., 2021b) 2021 FA+DA N/A CK+/Oulu-CASIA/MMI 97.85%/89.23%/75.32% CNN 10 fold+Hold out
Chen et al. (Chen et al., 2022b) 2022 FD+FA Atten Multi-PIE/KDEF 88.41%/89.04% CNN 5 fold
@@
Li et al. (Li et al., 2019b) 2018 FD+FA RES+Atten RAF-DB/AffectNet 85.07%/58.78% CNN 10 fold
Zhao et al. (Zhao et al., 2021a) 2021 FA+DA RES+Atten CAER/AffectNet/RAF-DB/SFEW 2.0 88.42%,/64.53%,/88.40%,/59.40% CNN Cross-validation
@@
Aouayeb et al. (Aouayeb et al., 2021) 2021 DA RES+Trans CK+/JAFFE/FER-2013/SFEW 2.0/RAF-DB 99.80%/92.92%/54.29%/87.22% CNN 10 fold + Hold out
Zheng et al. (Zheng et al., 2023) 2023 FA RES+Trans+Atten RAF-DB/AffectNet/FERPlus 92.05%/67.31%/90.83% CNN Hold out
@@
Sun et al. (Sun et al., 2023a) 2023 DA RES+Trans+Atten RAF-DB/AffectNet/FER2013 89.50%/65.66%/74.84% 2sCNN Hold out
@@
  • 1DA: Data augmentation; FD: Face detection; CE: Contrast enhancement; IN: Illumination normalization; FA: Face alignment; N/A: Not applicable;
  • 2Trans: Transformer; Atten: Attention; nsCNN: n-stream convolutional neural network; LOSO: Leave-one-subject-out; ROI: Region of interest; RES: Residual; GRU: Gated recurrent unit;

### v3

@@ Abstract.
Facial expressions convey human emotions and can be categorized into macro-expressions (MaEs) and micro-expressions (MiEs) based on duration and intensity. While MaEs are voluntary and easily recognized, MiEs are involuntary, rapid, and can reveal concealed emotions. The integration of facial expression analysis with Internet-of-Thing (IoT) systems has significant potential across diverse scenarios. IoT-enhanced MaE analysis enables real-time monitoring of patient emotions, facilitating improved mental health care in smart healthcare. Similarly, IoT-based MiE detection enhances surveillance accuracy and threat detection in smart security. Our work aims to provide a comprehensive overview of research progress in facial expression analysis and explores its potential integration with IoT systems. We discuss the distinctions between our work and existing surveys, elaborate on advancements in MaE and MiE analysis techniques across various learning paradigms, and examine their potential applications in IoT. We highlight challenges and future directions for the convergence of facial expression-based technologies and IoT systems, aiming to foster innovation in this domain. By presenting recent developments and practical applications, our work offers a systematic understanding of the ways of facial expression analysis to enhance IoT systems in healthcare, security, and beyond.
Facial expression analysis, Internet of Thing systems, macro- and micro-expressions
@@ 1. Introduction
Since its inception in 1999, Internet-of-Things (IoT) technology has seamlessly integrated into the fabric of modern society, driving key innovations in both civilian and industrial sectors. The widespread adoption of IoT technology generates skyrocketing amounts of daily data that were traditionally processed in clusters of cloud servers. Due to the ever-increasing privacy concerns, the information process is shifted from the cloud to the network edge. The low access latency and high data security of edge computing enable new application domains, such as emotion detection, traffic surveillance, and healthcare. Since facial expressions convey 55% of emotional messages during the exchange of moods, thoughts, feelings, and mental states, they have become a medium for sensing emotions in front of facility screens (Mehrabian and Wiener, 1967). Recent advancements in IoT technologies have driven the widespread application of facial expression analysis in practical scenarios. More specifically, IoT devices can first collect raw facial expression data for pre-processing. To fully leverage low access latency and high data security during the implementation of real-time privacy-preserving systems, the pre-processed features are then uploaded to edge devices for decision-making. Facial expressions can be divided into macro-expressions (MaEs) and micro-expressions (MiEs) based on their duration and intensity. In terms of duration, MaEs approximately last 0.5 to 4 seconds, while MiEs last less than 0.5 seconds (Shen et al., 2012). The short duration of MiEs conveys subtle emotional message, which makes emotion perception from MiEs a challenging task without specialized techniques and training. The differences between MaEs and MiEs are as follows.
@@
MaEs: can be perceived during regular interactions. The recognition accuracy of MaEs  can exceed 97% in a controlled laboratory due to the spontaneity and noise immunity. MaEs recognition can be applied in multiple scenarios, e.g., autonomous driving, psychological health assessment, and human-computer interaction; MiEs: are involuntary and rapid. Due to the short duration and subtle emotional cues, the detection of MiEs  can be challenging with naked eyes. It is reported that the average recognition accuracy of MiEs is around 47% after specialized training (Frank et al., 2009). Nevertheless, MiEs can reveal the true emotions of a person since they are instinctive and cannot be concealed. Numerous MiE applications in IoT systems are being explored, such as lie detection, criminal identification, and security control.
@@
The objective of facial expression analysis lies in detecting the emotions of humans from MaEs and MiEs. The research works on MaE analysis are mainly based on deep models that detect human emotions by leveraging the isolated static images and sequential dynamic images of MaEs. More specifically, the deep static MaE analysis performs well for two scenarios: (1) when several isolated images are available; and (2) when real-time expression recognition is required. The deep dynamic MaE analysis can benefit from the comprehensive understanding of temporal dynamics of MaEs. Since the emotion recognition of MiEs is challenging due to the short duration and subtle emotional cues, the research works on holographic MiE analysis consist of two parts, i.e., MiE spotting and MiE recognition. The MiE spotting aims at detecting MiEs within a given video, and the MiE recognition then uses the detected MiEs for MiE classification within a set of MiE categories. Different from the MaE analysis, the MiE analysis can provide valuable insights into potentially hidden human emotions.
@@ 2.2. Recent Surveys on MiE Analysis
Previous surveys on MiE analysis have comprehensively examined the field from various perspectives, e.g., MiE analysis, dataset creation, and psychological foundations of MiE development. However, critical aspects such as advanced MiE analysis and practical MiE applications remain underexplored. To address the aforementioned gaps, our work delves into state-of-the-art MiE analysis in Section 5. Additionally, we explore the transformative potential of MiE applications in the IoT system in Section 6.2.
Based on the previous discussion, we can summarize our unique contributions over the prior research on both MaE and MiE surveys as follows: 1). While existing surveys treat MaE and MiE in isolation, our work bridges this gap by providing a holistic framework that integrates both, addressing their complementary roles in facial expression analysis; 2). We propose a learning paradigm-based framework that unifies emerging techniques in deep MaE recognition and holographic MiE analysis that enable the systematic and insightful comparisons across these fields; 3). While much of the existing research emphasizes theoretical advancements in facial expression analysis, our work distinguishes itself by focusing on its practical deployment in real-world IoT scenarios, with a focus on potential applications and key implementation challenges. By unifying fundamental research, emerging technologies, and practical applications, our work offers a holistic framework that advances facial expression analysis within IoT systems. We envision that our work can foster continuous innovation by enhancing the development of emotion-aware applications across domains, e.g., smart healthcare and security.
@@
RaFD (Langner et al., 2010) 67 1 8040 images 256 × 256
@@
Oulu-CASIA (Zhao et al., 2011) 80 2
@@
EmotioNet (Fabian Benitez-Quiroz et al., 2016) N/A N/A 950,000 images Multiple
@@
AFEW (Dhall et al., 2017) N/A N/A 1809 sequences Multiple BEs N/A N/A N/A Movie P & S 2017
@@
Aff-Wild2 (Kollias and Zafeiriou, 2019) 554 N/A 564 sequences Multiple
@@
FERV39k (Wang et al., 2022e) N/A N/A 38,935 sequences 3840 × 2160 BEs N/A N/A N/A Internet P & S 2021
@@
  • 1 BEs: Basic expressions (seven); N/A: Not applicable; P: Posed; S: Spontaneous; NAF: Onset frame, Apex frame, Offset frame, respectively.
3. Datasets for Facial Expression Analysis
3.1. MaE Datasets
The recent MaE datasets including Japanese female facial expression (JAFFE) (Lyons et al., 1998), Karolinska directed emotional faces (KDEF) (Goeleven et al., 2008), MMI (Valstar et al., 2010; Pantic et al., 2005), Binghamton university 3D facial expression (BU-3DFE) (Yin et al., 2006), the Toronto face database (TFD) (Susskind et al., 2010), The Extended Cohn-Kanade Dataset (CK+) (Lucey et al., 2010), Radboud faces database (RaFD) (Langner et al., 2010), MUG (Aifanti et al., 2010), Multi-PIE (Gross et al., 2010), static facial expressions in the wild (SFEW 2.0) (Dhall et al., 2011), Oulu-CASIA (Zhao et al., 2011), facial expression recognition 2013 (FER-2013) (Goodfellow et al., 2013), Indian spontaneous expression database (ISED) (Happy et al., 2017), EmotioNet (Fabian Benitez-Quiroz et al., 2016), wild affective faces database (RAF-DB) (Li et al., 2017a), AffectNet (Mollahosseini et al., 2019), the acted facial expressions in the wild (AFEW) (Dhall et al., 2017), expression in-the-wild (ExpW) (Zhang et al., 2018b), Aff-Wild2 (Kollias and Zafeiriou, 2019), context-aware emotion recognition (CAER) (Lee et al., 2019), dynamic facial expression in the wild (DFEW) (Jiang et al., 2020), FERV39k (Wang et al., 2022e), Padova emotional dataset of facial expressions (PEDFE) (Miolla et al., 2023), CalD3r + MenD3s (Ulrich et al., 2024). Table 1 illustrates the detailed information of the MaE datasets.
@@ 3.2. MiE Datasets
In recent years, multiple datasets have been built for further development of MiE analysis. Initially, researchers collected the MiE databases (e.g., Polikvsky’s database (Polikovsky et al., 2009) and USF-HD (Shreve et al., 2011)) by asking the participants to pose or mimic facial expressions. These posed datasets supported the preliminary studies of MiE, but were not used for current analysis due to the properties of private datasets and not capturing the nature of MiE. Later, more spontaneous datasets, including the spontaneous micro-expression corpus (SMIC) (Li et al., 2013), Chinese Academy of Sciences micro-expression (CASME) (Yan et al., 2013), Chinese Academy of Sciences micro-expression II (CASME II) (Yan et al., 2014), Chinese Academy of Sciences macro-expressions and micro-expressions (CAS(ME)2) (Qu et al., 2017), spontaneous actions and micro-movements (SAMM) (Davison et al., 2016), micro-expression videos in the wild (MEVIEW) (Husák et al., 2017), CAS(ME)3 (Li et al., 2022a), MMEW (Ben et al., 2021), and 4D spontaneous micro expression database (4DME) (Li et al., 2022b), were utilized in MiE research. The information on the spontaneous MiE datasets is detailed in Table 2.
@@
Figure 1. Samples of current MiE datasets. These are as follows: (a) SMIC (Li et al., 2013), (b) CASME (Yan et al., 2013), (c) CASME II (Yan et al., 2014), (d) CAS(ME)2 (Qu et al., 2017), (e) SAMM (Davison et al., 2016), (f) MEVIEW (Husák et al., 2017), (g) CAS(ME)3 (Li et al., 2022a), (h) MMEW (Ben et al., 2021), and (i) 4DME (Li et al., 2022b).
@@
3.3. Challenges in Practical Applications of Facial Expression Datasets
Despite the significant progress in developing MaE and MiE datasets, several challenges persist when applying them to real-world scenarios. Four key challenges that hinder the practical application of facial expression datasets are discussed as follows. Class Imbalance and Label Inconsistency: Overrepresentation of common expressions (e.g., happiness and sadness) versus underrepresentation rare ones (e.g., contempt, fear, and disgust). Besides, inconsistencies in expression classification exist for both MaE and MiE datasets. Collection Complexity: Facial expression data collection poses substantial challenges. Spontaneous expressions are inherently more difficult to elicit and capture than posed ones, making the collection of spontaneous datasets more difficult. Ensuring demographic diversity—such as including participants across ethnicities and age groups—further complicates the data collection process. Additionally, MiE datasets require high-frame-rate recordings to capture subtle and transient facial movements, imposing strict demands on both recording equipment and experimental design. Annotation Complexity: MiE labeling is very labor-intensive and subject to inter-annotator variability, even among trained experts, which further reduces reliability. Meanwhile, the large scale of MaE datasets demands automated annotation strategies to for data labeling. Domain Shift: Domain Shift arises both across datasets within the same task (e.g., MaE and MiE) and between training data and real-world deployment scenarios. Variations in collection protocols, environmental conditions, and label annotation often lead to inconsistent model performance across datasets. Moreover, the mismatch between lab-controlled data and in-the-wild expressions further hinders the robustness of real-world applications.
@@ 3.3. Challenges in Practical Applications of Facial Expression Datasets
Pons et.al . (Pons and Masip, 2018a) 2018 IN+FA N/A SFEW 2.0 42.9% CNN LOSO
Saurav et al. (Saurav et al., 2022) 2022 IN+FA+CE N/A FER2013/RAF-DB/CK+ 72.77%/86.07%/98.54% CNN LOSO
Shao et al. (Shao and Qian, 2019) 2019 FD+FA+DA RES CK+/BU-3DFE/FER2013 95.29%/86.50%/71.14% CNN 10 fold
@@
Wadhawan et al. (Wadhawan and Gandhi, 2023) 2023 FD+CE N/A CK+/JAFFE 97.31%/97.14% CNN 10 fold+LOSO
@@
Atabansi et al. (Atabansi et al., 2021) 2021 FD+DA N/A Oulu-CASIA NIR 98.11% CNN Hold out
Multi-task Learning Pons et al. (Pons and Masip, 2018b) 2018 DA RES SFEW 2.0 45.9% CNN Hold out
@@
Yu et al. (Yu and Xu, 2022) 2022 DA+FA RES+Atten RAF-DB/SFEW 2.0/CK+/Oulu-CASIA 90.36%/45.78%/98.33%/87.32% CNN 10 fold+LOSO
Liu et al. (Liu et al., 2024) 2023 N/A RES+Trans AffectNet/Aff-Wild2 65.80%/64.13% CNN Hold out
Yu et al. (Yu et al., 2020) 2020 FA+ROI N/A CK+/Oulu-CASIA 99.08%/90.40% GRU 10 fold
Xiao et al. (Xiao et al., 2023) 2023 FD N/A CK+/MMI/RAF-DB 99.07%/84.62%/87.52% CNN 5 fold+Hold out
Zhao et al. (Zhao et al., 2021a) 2021 FA+DA N/A CK+/Oulu-CASIA/MMI 97.85%/89.23%/75.32% CNN 10 fold+Hold out
Chen et al. (Chen et al., 2022a) 2022 FD+FA Atten Multi-PIE/KDEF 88.41%/89.04% CNN 5 fold
@@
Li et al. (Li et al., 2019c) 2018 FD+FA RES+Atten RAF-DB/AffectNet 85.07%/58.78% CNN 10 fold
Zhao et al. (Zhao et al., 2021c) 2021 FA+DA RES+Atten CAER/AffectNet/RAF-DB/SFEW 2.0 88.42%/64.53%/88.40%/59.40% CNN Cross-validation
@@
Aouayeb et al. (Aouayeb et al., 2021) 2021 DA RES+Trans CK+/JAFFE/SFEW 2.0/RAF-DB 99.80%/92.92%/54.29%/87.22% CNN 10 fold + Hold out
Zheng et al. (Zheng et al., 2023) 2023 FA RES+Trans+Atten RAF-DB/AffectNet/FERPlus 92.05%/67.31%/91.62% CNN Hold out
@@
Sun et al. (Sun et al., 2023b) 2023 DA RES+Trans+Atten RAF-DB/AffectNet/FER2013 89.50%/65.66%/74.84% 2sCNN Hold out
@@
  • 1DA: Data augmentation; FD: Face detection; CE: Contrast enhancement; IN: Illumination normalization; FA: Face alignment; N/A: Not applicable; Trans: Transformer; Atten: Attention; nsCNN: n-stream convolutional neural network; LOSO: Leave-one-subject-out; ROI: Region of interest; RES: Residual; GRU: Gated recurrent unit.